An elevator is moving up at a constant velocity of 2.50 m/s the man has a mass of 85.0 kg how large is the force exerts on the man


1
Expert's answer
2022-04-11T12:07:36-0400
"F=m(g+a)\\
=85.0*(9.81+0)=834\:\rm N"